# spacing=nonfrench

### v1

@@ 6.1 Architecture of GenAI browser assistants
This section discusses our findings related to RQ1, providing a deeper understanding on how GenAI browser assistants are designed. We qualitatively analyze network traffic activity to learn the following architectural differences across 10 assistants: (1) backend model, (2) response architecture, (3) context restrictions, and (4) response variability.
@@
Backend model. Since GenAI browser assistants function as wrappers on top of the open source LLM models, we observe that they provide support for most of the popular open-source models – OpenAI’s chatGPT [36], Google’s Gemini [16], Anthropic’s Claude AI [4], and Meta’s Llama [26] as shown in Table 1. For each extension, the request to fetch the response also contained the model information. We notice gpt-4o-mini to be the most popularly used default model due to its cost-efficiency trade-off for assistant developers as well as for users over other assistants. Some extensions used different models for different user activities. For example, ChatGPT for Google (CFG) used gpt-4o-mini for Google search triggered automatic response generation while it used gpt-3.5-turbo when a user explicitly opened a sidebar chat to ask some question. Sider, Perplexity, and Harpa use custom LLM models, making it difficult to understand how its design differs from other open-source models.
@@
Full Webpage : Page text, title, location, hyperlinks. Server-fetch Webpage : Page title, location, server-fetched file’s upload location. Plain Webpage : Page text, title, location. Partial Webpage : Partial content or missing details. Response legend: ✔: Response with Relevant Details. ✔: Missing some details in Response. : Response restricted. ✘: No response generated.
@@
Category WebPage Sider Monica CFG Merlin MaxAI Perplexity Harpa Wiseone TinaMind Copilot Public Spaces News Platforms cnn.com Open Forums reddit.com Informative Articles wikipedia.org E-commerce Website amazon.com Sports Websites espn.com Travel Platforms expedia.com User-generated Media youtube.com Kids Website nickjr.com Misinformation Website infowars.com Violence Material guns.com Private Spaces Health Portal university health portal Email Account mail.google.com Social Media Platform facebook.com Adult Content pornhub.com Online Streaming Service netflix.com Government Website irs.gov Dating Service tinder.com Financial Service chase.com Educational Platform canvas.instructure.com Messaging Platform slack.com
Response architecture. Browser assistants may contact open-source GenAI models to generate responses either from client-side or server-side. Figure 1 depicts the most common case we observed in our study, where 9 out of 10 extensions operate server-side. This means that first, the user’s selected model (or default in case of no selection) is shared along with the user’s query and other metadata to the browser assistant’s server. Based on the choice of model, assistant’s server would then call API of the respective LLM model to direct the user’s request to. The personalized response is finally displayed to the user. In case of CFG, response generation is initiated from the client-side (see Table 1). To use CFG browser assistant, the user needs to link their OpenAI’s chatgpt.com account with CFG. As a result, all queries from interaction with CFG are directly shared with chatgpt.com to fetch the response. We also analyze if browser assistants always monitor user’s activities (passively) or if they perform monitoring only when user explicitly interacts with the assistant (actively). Here we refer to activities that do not exist as features in the browser assistant, such as scrolling, clicks, navigation. etc. We found it to be always “active” for all 10 extensions, suggesting that they do not passively monitor user’s non-extension related activities. In regards to extension-related activities like search, we found that some assistants such as Sider and Copilot automatically invoke response generation when a user submits a search query in the search engine. 5/10 extensions require manual invocation of the assistant to obtain a response while remaining 3 work in a mixed fashion.
Context restrictions. Next, we compare the interdependence of context used by the browser assistant under different scenarios. We look at two sets of context – sidebar response versus popup sidechat and context across page navigations. In the former case, during our experiments, we analyse dependency of context from the response generated by the assistant in the empty sidebar space of Google search results page with the context of the sidebar chat that pops up upon clicking the extension icon. We did a simple experiment, where we asked in the Google search ‘I like apples’ (Q1) and ensured that the sidebar response shows up. Next, we opened the sidebar popup chat and asked ‘Do I like apples?’ (Q2) – if it responds ‘Yes’ then that suggests the two contexts are interpreted as dependent, otherwise independent. We observed that 4 assistants – ChatGPT for Google, Max AI, TinaMind, and Copilot maintained the same context while the remaining 6 maintained separate contexts, suggesting stronger privacy. In the second case, we ask Q1 in the sidebar popup chat when visiting some webpage; next we navigate to a different webpage and ask Q2 to understand if it preserved the context or not. We observe that context was not reset only for 2/10 assistants – ChatGPT for Google and WiseOne, suggesting their capability to remember contexts across multiple websites as user browses through the web. We found Perplexity to be the most private assistant as it explicitly displayed the message stating ‘I do not have the ability to recall previous interactions or questions. Each session or question is treated independently for privacy reasons.’. We further analyze what data forms the necessary context for different assistants in more detail in Section 6.2
Response Variability. It is important to understand that despite different browser assistants using the same model to perform the user’s query, the outputs returned to the users could be completely different. Several factors play a role in this. We discuss two of the most important ones below – Temperature and System prompts.
Temperature. First, as aforementioned, output of generative models is probabilistic and is influenced by the temperature parameter. For example, OpenAI’s ChatGPT provides capability to set temperature in the range 0 to 1 where lower values provide more deterministic and focused outputs while higher values generate more creative or random responses [35]. Different browser assistants may set and use different values of temperature, resulting in differences in outputs even when they use the same model. Moreover, even the same browser assistant may dynamically set temperature for different user queries. For example, when asked to a browser assistant: “What do you know about me?” – it may query OpenAI’s API with a higher temperature value to not generate an extremely precise (or invasive) output and rather keep it open-ended. However, if the same user asks the same browser assistant: “What is the distance between New York and Washington D.C.?”, the assistant may use a smaller value of temperature to provide a deterministic output.
@@
System Prompts. Second, system prompts are provided to the LLM models along with the user query for various reasons such as to enhance task-specific accuracy, provide contextual guidance, prevent unwanted output, ensure output formatting, and mitigate vulnerabilities such as jailbreaking [48]. During the network traffic analysis, we discovered various system prompts being empoloyed by different browser assistants as listed in Section 8.3 in appendix. We observe some assistants to use a tighter system prompt to answer a user’s query more accurately than others. For example, Sider’s system prompt to answer a question simply states to “use simple and clear language” whereas Harpa’s system prompt explicity states to “NEVER fabricate, infer, or guess information. Do not hallucinate links. Be to the point”. In the same prompt, it can be seen that Harpa includes $\{\{user\_info\}\}$ as a context along with user’s query to generate a response, suggesting that it shares user details with open-source LLM models freely with each query. Moreover, Harpa’s system prompt states “Please ignore all previous instructions” to prevent any jail-breaking attempts through cleverly-crafted user queries. TinaMind’s system prompts are observed to associate language preference, day, date, timestamp, and timezone information with each query in order to interpret the user’s location-dependent queries as per the applicable region to produce a more relevant response.
@@ 6.2 User tracking by GenAI browser assistants
To systematically audit user tracking by different browser assistants, in Section 6.2.1, we first perform analysis of implicit collection and sharing of user’s data as user visits public and private spaces online. We refer to it as ‘implicit’ as the user does not want to actually share their personal information with the GenAI assistant. However, the user may inadvertently let the assistant access it. Here, we analyse tracking of user data in context of the webpage content that the user is visiting. Next, in Section 6.2.2, we evaluate explicit data collection and sharing when using different functionalities offered by assistants, where we use the prompting framework to explicitly leak user attributes and understand tracking of all kinds of data associated with the user.
@@ 6.2.1 Implicit collection and sharing of user data
In this section, we discuss the results related to implicit user tracking to see if and when do browser assistants collect data. We followed the methodology described in Section 5.3 to visit different websites in public and private spaces, summarize the page content, and ask a follow-up question. Table 2 showcases results from running a total of 200 experiments across 20 websites and 10 assistants.
@@
As part of the page data, we consider page title, page text, page location and any embedded links on the page. The cells in red suggests most egregious data collection, where complete webpage details were collected, while cells in green correspond to inefficient collection of pages (e.g., with missing details). We observe that Harpa collects full DOM in all 20 online spaces, while TinaMind is unable to extract full webpage details across all the experiments. This is likely due to two reasons – one, TinaMind only focuses at capturing data that is in the viewport of the user, failing to capture rest of the page and two, its page extraction mechanism is less aggressive. The latter inference is based on the fact that it performed incomplete extraction even from the user’s viewport content, often failing to reproduce the numbers present on the webpage in its responses.
On their Chrome webstore page, Harpa clearly states “we do not collect or sell userdata” – contradictorily, we observe it to collect 100% of health records, student academic data, as well as user’s personal messages on messaging platforms. As mentioned in Section 6.1, Harpa’s system prompt also included the user’s name and location in plain text, for example: “My name is John Gabb. I am in London, United Kingdom of Great Britain and Northern Ireland. Please answer in English…” This demonstrates clear violation of their own privacy practices.
@@
Browser assistants can easily employ practices that can allow them to determine if a page contains sensitive or private data or not. For instance, when summarize feature is used with Perplexity, it shares the page URL with its own server and performs a server-side fetch of the webpage. This is inferred based on the file upload location received in response to the shared page location. Therefore, if the webpage belongs to a user’s private authenticated space, Perplexity’s server will never have access to the user’s personal data. However, it will still be able to provide responses for public spaces. In contrast, Wiseone employs a different approach by obtaining explicit user consent for each new website where the user wishes to utilize its functionalities.
@@
Merlin, MaxAI, ChatGPT for Google and Monica were all able to extract webpage contents for all 20 scenarios. Merlin was discovered to be the only assistant that recorded even the contents of the forms on webpages as opposed to all other browser assistants – that did not collect form data. For example, it was able to collect the user’s “Social Security Number (SSN)” entered in a form on IRS refund portal. Email addresses as well as the full email thread were also collected by these assistants. One of the most shocking findings was that GenAI browser assistants were freely able to collect and share data to their own servers on authenticated health portals. They were able to answer follow-up questions ranging from patient details to entire medical history. Collection of PHI without appropriate user consent is in clear violation of HIPAA [14]. Moreover, student’s academic records including assessment scores, exam performances, overall grades – were all collected and shared with browser assistant’s servers demonstrating violation of FERPA [34] that aims to protect these attributes for a student.
@@
Overall, it can be observed that responses with missing details are more prominent in the lower half of the table suggesting that private spaces are handled with care by most extensions. However, this is insufficient since they are still collecting partial data that is private to the users. Adult content, online streaming services, dating platforms, and financial services – were unanimously responded inefficiently by all the assistants. Titles displayed on Netflix homepage, preferences set on Tinder and transaction amounts as well as last four digits of card number on Chase were either inaccurately output or were missing from the output of different assistants. This is because we observed that the assistants missed capturing shadow elements from DOM of the webpage. On the brighter side, Perplexity and Wiseone actively suppressed response generation on adult pages due to explicit content stating “I apologize, but I cannot provide information about or analyze pornographic content. This type of material is not appropriate for me to discuss. Perhaps I could assist you with a different, non-explicit topic instead?”. However, both of them still recommended links to pornographic content in follow-up suggestions, showing inefficiencies in their model.
@@
6.2.2 Explicit collection and sharing of user data
Having looked at what online spaces are vulnerable to data collection, we now focus on understanding what attributes are explicitly collected about the user, device, or page when different features provided by the browser assistant are used by the user – namely, search, browse and summarize as shown in Table 3. We look at the requests, payloads, and headers to identify different attributes that are collected and shared with either first-party server of the browser assistant, third-party servers, both or none during different scenarios.
@@ 6.2.2 Explicit collection and sharing of user data
User’s Data. Now, we look at collection and sharing of user’s search query, chat details, and user-related identifiers with different endpoints. Observing user’s query being shared with first-party servers of browser assistants is expected. However, surprisingly, we found Merlin’s background service worker to include Google Analytics tracking script. As a result, user’s raw query was also shared with google-analytics.com endpoint. In case of CFG, we observed the query being shared with chatgpt.com to fetch a response to the user’s prompt. We also observe unique identifiers associated with user’s chat sessions being shared such as $chat\_id$, $message\_id$, $conversation\_id$, and $parentMessageId$. Such chat identifiers were shared with all first-party servers, except Perplexity and Wiseone. In case of TinaMind, chat identifiers were shared with analytics.google.com along with the user identifiers. However, Sider and Merlin were observed sharing similar user identifiers with google-analytics.com. An important distinction is that sharing data with google-analytics.com allows tracking the user for the purpose of analytics. However, sharing data with analytics.google.com allows joining user’s identity across Google’s domain google.com with shared cookies. Browser assistant developers can create custom audiences based on the query terms or chat identifiers to (re-)target users with ads across Google properties such as mail.google.com for instance. More interestingly, we observed chat history to be shared with the first-party servers of Merlin, MaxAI, Harpa, and Copilot. Harpa and Copilot maintained the entire chat history of the user since their first conversation in the background service worker’s IndexedDB storage and the full history was shared with every new query to provide complete context. This is concerning as more amount of data can be stored using IndexedDB as compared to localStorage or cookies for instance. Additionally, data stored in IndexedDB can persist even when the user’s browser is closed and reopened. Perplexity stored states using what it referred to as rwToken in localStorage. In terms of cookies, we observe many extensions setting first-party cookies like _fbp (Facebook), _ga (Google analytics), _clk (Clarity), etc. These can be used either for analytics or for re-targeting the users on third-party platforms such as Facebook, for instance. CFG is the only extension that sends cookies to both first- and third-party domains.
@@
Thus, third-party data collection and sharing is more concerning than first-party as it allows linking and targeting of user across multiple websites. Figure 3 summarizes all third-party sharing observed by us. Moreover, first-party data collection and sharing suggests potential of profiling and personalization as we discuss in the next section.

### v2

@@ 6.1 Architecture of GenAI browser assistants
This section discusses our findings related to RQ1, providing a deeper understanding of how GenAI browser assistants are designed. We qualitatively analyze network traffic activity to learn the following architectural differences across 9 assistants: (1) backend model, (2) response architecture, (3) context restrictions, and (4) response variability.
@@
Backend model. Since GenAI browser assistants function as wrappers on top of the proprietary LLM models, we observe that they provide support for most of the popular models – OpenAI’s chatGPT [36], Google’s Gemini [15], Anthropic’s Claude AI [3], and Meta’s Llama [23] as shown in Table 1. For each extension, the request to fetch the response also contained the model information. We observed gpt-4o-mini to be the most popularly used default model due to its cost-efficiency trade-off for assistant developers as well as for users over other models. Some extensions used different models for different user activities. For example, ChatGPT for Google (CFG) used gpt-4o-mini for Google-search-triggered automatic response generation while it used gpt-3.5-turbo when a user explicitly opened a sidebar chat to ask some question. Sider, Perplexity, and Harpa use custom LLM models, making it difficult to determine how their designs differ from commonly used proprietary models.
Response architecture. Browser assistants may contact popular proprietary GenAI models to generate responses either from client-side or server-side. Figure 1 depicts the most common case we observed in our study, where 8 out of 9 extensions operate server-side. This means that first, the user’s selected model (or default in case of no selection) is shared along with the user’s query and other metadata to the browser assistant’s server. Based on the selected model, the assistant’s server calls the corresponding LLM API to direct the user’s request to. The personalized response is finally displayed to the user. In case of CFG, response generation is initiated from the client-side (see Table 1). To use CFG browser assistant, the user needs to link their OpenAI’s chatgpt.com account with CFG. As a result, all queries from interaction with CFG are directly shared with chatgpt.com to fetch the response. We also analyze if browser assistants always monitor user’s activities (passively) or if they perform monitoring only when user explicitly interacts with the assistant (actively). Here, we refer to browser activities such as scrolling, clicks, navigation, etc. that is not tied to explicit assistant features. We found it to be always “active” for all 9 extensions, suggesting that they do not passively monitor users’ non-extension-related activities. In regards to extension-related activities like search, we found that some assistants such as Sider and Copilot auto-invoke response generation when a user submits a search query in the search engine. 4/9 extensions require manual invocation of the assistant to obtain a response while remaining 3 work in a mixed fashion.
Context restrictions. Next, we compare the interdependence of context used by the browser assistant under different scenarios. We look at two sets of context – sidebar response versus popup sidechat and context across page navigations. In the former case, during our experiments, we analyze dependency of context from the response generated by the assistant in the empty sidebar space of Google search results page with the context of the sidebar chat that pops up upon clicking the extension icon. We conducted a simple experiment in which we searched ‘I like apples’ on Google Search (Q1) and ensured that the sidebar response shows up. Next, we opened the sidebar popup chat and asked ‘Do I like apples?’ (Q2) – if it responds ‘Yes’ then that suggests the two contexts are dependent, otherwise independent. We observed that 4 assistants – ChatGPT for Google, Max AI, TinaMind, and Copilot maintained the same context while the remaining 5 maintained separate contexts, suggesting stronger privacy. In the second case, we ask Q1 in the sidebar popup chat while visiting a specific webpage; next we navigate to a different webpage and ask Q2 to determine whether the assistant preserved context. We observe that context was not reset only for one assistant – ChatGPT for Google, suggesting its capability to remember contexts across multiple websites as user browses through the web. We found Perplexity to be the most private assistant as it explicitly displayed the message stating ‘I do not have the ability to recall previous interactions or questions. Each session or question is treated independently for privacy reasons.’. We further analyze what types of data form the necessary context for different assistants in more detail in Section 6.2.
Response Variability. It is important to understand that even when different browser assistants use the same model to process a user’s query, the output may vary significantly. Several factors play a role in this. We discuss two of the most important ones below: Temperature and System prompts.
Temperature. First, as aforementioned, output of generative models is probabilistic and is influenced by the temperature parameter. For example, OpenAI’s ChatGPT provides the capability to set temperature in the range of 0 to 1, where lower values provide more deterministic and focused outputs while higher values generate more creative or random responses [35]. Different browser assistants may set and use different values of temperature, resulting in differences in outputs even when they use the same model. Moreover, even the same browser assistant may dynamically adjust temperature based on the nature of the user query. For example, when asked to a browser assistant: “What do you know about me?” – it may query OpenAI’s API with a higher temperature value to avoid generating a precise response and rather keep it open-ended. However, if the same user asks the same browser assistant: “What is the distance between New York and Washington D.C.?”, the assistant may use a smaller value of temperature to provide a deterministic output.
System Prompts. Second, system prompts are provided to the LLM models along with the user query for various reasons such as to enhance task-specific accuracy, provide contextual guidance, prevent unwanted output, ensure output formatting, and mitigate vulnerabilities such as jailbreaking [45]. During the network traffic analysis, we discovered various system prompts being employed by different browser assistants as listed in Section 8.2 of the appendix. We observe that some assistants use stricter system prompts to answer a user’s query more accurately than others. For example, Sider’s system prompt to answer a question simply states to “use simple and clear language” whereas Harpa’s system prompt explicitly states to “NEVER fabricate, infer, or guess information. Do not hallucinate links. Be to the point”. In the same prompt, it can be seen that Harpa includes $\{\{user\_info\}\}$ as context
@@
Full Webpage: Page text, title, location, hyperlinks. Server-fetch Webpage: Page title, location, server-fetched file’s upload location. Plain Webpage: Page text, title, location. Partial Webpage: Partial content or missing details. Response legend:
✔: Response with relevant details. ✔: Missing some details in response : Response restricted. ✘: No response generated.
@@
Category WebPage Sider Monica CFG Merlin MaxAI Perplexity Harpa TinaMind Copilot Public Spaces News Platforms cnn.com Open Forums reddit.com Informative Articles wikipedia.org E-commerce Website amazon.com Sports Websites espn.com Travel Platforms expedia.com User-generated Media youtube.com Kids Website nickjr.com Misinformation Website infowars.com Violence Material guns.com Private Spaces Health Portal hem.ucdavis.edu Email Account mail.google.com Social Media Platform facebook.com Adult Content pornhub.com Online Streaming Service netflix.com Government Website irs.gov Dating Service tinder.com Financial Service chase.com Educational Platform canvas.instructure.com Messaging Platform slack.com
@@
along with the user’s query to generate a response, suggesting that it shares user details with proprietary LLM models freely with each query. Moreover, Harpa’s system prompt states “Please ignore all previous instructions” to prevent any jail-breaking attempts through cleverly-crafted user queries. TinaMind’s system prompts include user-specific metadata such as language preference, timestamp, and timezone details with each query in order to interpret the user’s location-dependent queries as per the applicable region to produce a more relevant response.
@@ 6.2 User tracking by GenAI browser assistants
To systematically audit user tracking by different browser assistants, in Section 6.2.1, we first perform analysis of implicit collection and sharing of user’s data as user visits public and private spaces online. We refer to it as ‘implicit’ as the user does not actually intend to share their personal information with the GenAI assistant. However, the user may inadvertently let the assistant access it. Here, we analyze tracking of user data in context of the webpage content that the user is visiting. Next, in Section 6.2.2, we evaluate explicit data collection and sharing when using different functionalities offered by assistants, where we use the prompting framework to explicitly leak user attributes and understand tracking of all kinds of data associated with the user. Lastly, in Section 6.2.3, we qualitatively compare the observed data collection and tracking practices against the privacy policy of the assistants.
@@ 6.2.1 Implicit collection and sharing of user data
In this section, we discuss the results related to implicit user tracking to see if and when do browser assistants collect data. We followed the methodology described in Section 5.3 to visit different websites in public and private spaces, summarize the page content, and ask a follow-up question. Table 2 showcases results from running a total of 180 experiments across 20 websites and 9 assistants.
@@
As part of the page data, we consider page title, page text, page location and any embedded links on the page. The cells in orange red suggests most egregious data collection, where complete webpage details were collected, while cells in green correspond to inefficient collection of pages (e.g., with missing details). We observe that Harpa collects full DOM in all 20 online spaces, while TinaMind is unable to extract full webpage details across all the experiments. This is likely due to two reasons – one, TinaMind only focuses at capturing data that is in the viewport of the user, failing to capture rest of the page and two, its page extraction mechanism is less aggressive. The latter inference is based on the fact that it performed incomplete extraction even from the user’s viewport content, often failing to reproduce the numerical content present on the webpage in its responses.
@@
Browser assistants can easily employ practices that can allow them to determine if a page contains sensitive or private data or not. For instance, when summarize feature is used with Perplexity, it shares the page URL with its own server and performs a server-side fetch of the webpage. This is inferred based on the file upload location received in response to the shared page location. Therefore, if the webpage belongs to a user’s private authenticated space, Perplexity’s server will never have access to the user’s personal data. However, it will still be able to provide responses for public spaces.
@@
Merlin, MaxAI, ChatGPT for Google and Monica were all able to extract webpage contents for all 20 scenarios. Merlin was discovered to be the only assistant that recorded even the contents of the forms on webpages as opposed to all other browser assistants – that did not collect form data. For example, it was able to collect the user’s Social Security Number (SSN) entered in a form on IRS refund portal. Email addresses as well as the full email thread were also collected by these assistants. One of the most surprising findings was that GenAI browser assistants collected and shared data to their own servers on authenticated patient portals. They were able to answer follow-up questions ranging from patient details to their medical history. Collection of such sensitive information without opt-in user consent is a potential violation of HIPAA [32]. Moreover, student academic records including assessment scores, exam performances, overall grades – were all collected and shared with browser assistant’s servers, which is a potential violation of FERPA [33].
@@
Overall, it can be observed that responses with missing details are more prominent in the lower half of the table suggesting that private spaces are handled with care by most extensions. However, this is insufficient since they are still collecting partial data that is private to the users. Adult content, online streaming services, dating platforms, and financial services – were unanimously responded inefficiently by all the assistants. Titles displayed on Netflix homepage, preferences set on Tinder and transaction amounts as well as last four digits of card number on Chase were either inaccurately output or were missing from the output of different assistants. This is because we observed that the assistants missed capturing shadow elements from DOM of the webpage. On the brighter side, Perplexity actively suppressed response generation on adult pages due to explicit content stating “I apologize, but I cannot provide information about or analyze pornographic content. This type of material is not appropriate for me to discuss. Perhaps I could assist you with a different, non-explicit topic instead?”. However, Perplexity still recommended links to pornographic content in follow-up suggestions, showing inefficiencies in their model.
6.2.2 Explicit collection and sharing of user data
Having looked at what online spaces are vulnerable to data collection, we now focus on understanding what attributes are explicitly collected about the user, device, or page when different features provided by the browser assistant are used by the user – namely, search, browse and summarize as shown in Table 3.
@@ 6.2.2 Explicit collection and sharing of user data
We look at the requests, payloads, and headers to identify different attributes that are collected and shared with either first-party server of the browser assistant, third-party servers, both or none during different scenarios.
@@
User Data. Now, we look at collection and sharing of user’s search query, chat details, and user-related identifiers with different endpoints. Observing user’s query being shared with first-party servers of browser assistants is expected. However, surprisingly, we found Merlin’s background service worker to include Google Analytics tracking script. As a result, user’s raw query was also shared with google-analytics.com endpoint. In case of CFG, we observed the query being shared with chatgpt.com to fetch a response to the user’s prompt. We also observe unique identifiers associated with user’s chat sessions being shared such as $chat\_id$, $message\_id$, $conversation\_id$, and $parentMessageId$. Such chat identifiers were shared with all first-party servers, except Perplexity. In case of TinaMind, chat identifiers were shared with analytics.google.com along with the user identifiers. However, Sider and Merlin were observed sharing similar user identifiers with google-analytics.com. An important distinction is that sharing data with google-analytics.com allows tracking a user for the purpose of analytics. However, sharing data with analytics.google.com allows joining user’s identity across Google’s domain google.com with shared cookies. Browser assistant developers can create custom audiences based on the query terms or chat identifiers to (re-)target users with ads across Google properties such as mail.google.com for instance. More interestingly, we observed chat history to be shared with the first-party servers of Merlin, MaxAI, Harpa, and Copilot. Harpa and Copilot maintained the entire chat history of the user since their first conversation in the background service worker’s IndexedDB storage and the full history was shared with every new query to provide complete context. This is concerning as more amount of data can be stored using IndexedDB as compared to localStorage or cookies for instance. Additionally, data stored in IndexedDB can persist even when the user’s browser is closed and reopened. Perplexity stored states using what it referred to as rwToken in localStorage. We also observe many extensions setting first-party cookies like _fbp (Facebook), _ga (Google analytics), _clk (Clarity), etc. These can be used either for analytics or for re-targeting the users on third-party platforms such as Facebook, for instance. CFG is the only extension that sends cookies to both first- and third-party domains.
@@
Thus, third-party data collection and sharing is more concerning than first-party as it allows linking and targeting of a user across websites. Figure 3 summarizes all third-party sharing observed by us. Moreover, first-party data collection and sharing suggests potential of profiling and personalization that we discuss in the next section.
Figure 3: Sankey diagram depicting third-party flows of different assistants. Thickness represents the # of requests.
6.2.3 Privacy policies of Gen AI browser assistants
To understand the disclosures about data collection and sharing practices of the browser assistants, we perform a qualitative analysis of their privacy policies and compare them with observed behaviors. Overall, we find widespread non-compliance and incomplete disclosures across the policies. All assistants either directly state or hint towards not collecting any personal or sensitive user information. However, none of them explicitly address the handling of webpage content that may contain such information, especially in private spaces. Consequently, the policies do not address how LLMs process user’s private data. This omission raises serious concerns about compliance and transparency, as many of these assistants collect health data and academic records, potentially violating HIPAA and FERPA. In addition, several assistants collect other personal information such as emails and private messages.
Next, we address privacy policies of individual assistants. Although Sider states that “We collect information about how you use the Services…”, it does not explicitly disclose collection of page content, chat IDs, and user queries. It does, however, acknowledge sharing user identifiers with third parties, consistent with our findings (“We may share your personal information with third-party service providers…”). Monica and Merlin discloses collection of personal details (e.g., name, email), IP address and cookies, but does not explicitly acknowledge about collecting Google search results (or form data in case of Merlin) or more generally, page content. In fact, Monica claims in its privacy policy that it “operates without intruding into your personal activities online. We collect no information on the sites you visit or the content you engage with.” MaxAI makes the same claim. Contrary to those statements, we observe Monica collecting webpage data across nearly all public and private sites we tested. Similarly, MaxAI shares page URLs, referrers, timestamps, and user details with both its own servers and Mixpanel. It also sends the full page content and chat history. These practices are unacknowledged in their privacy policies, which explicitly promise not to collect such data. Merlin acknowledges sharing of data with third parties “for authorized purposes”, without actually listing details on what constitutes an authorized purpose.
TinaMind’s privacy policy states “We do not share your personal information with third parties except as necessary to provide our products or services to you”. However, we observe chat and user identifiers being shared with analytics.google.com. Similarly, on their Chrome Web Store page, Harpa clearly states “we do not collect or sell userdata” – contradictorily, we observe it to collect user data as part of webpage data extracted in private spaces. Its privacy policy also states that it uses “Personal Data” (e.g., name and email) to “contact you with newsletters, marketing or promotional materials”. However, as mentioned in Section 6.1, we observe Harpa’s LLM system prompt to also include the user’s name and location in plain text, for example: “My name is John Doe. I am in London …” This use case is not addressed in its privacy policy.
In contrast, Copilot, CFG, and Perplexity are aligned with their stated policies. For example, Copilot discloses the collection of usage, device, and profile data, which aligns with our observations–though it omits mention of cookie usage. CFG’s privacy policy is the same as Monica’s policy, with the only difference being that it acknowledges that information could be shared with third parties, which matches our findings. Perplexity’s privacy policy discloses all data collection and sharing practices that were observed by us.